OpenStudy (anonymous):

Choose the correct graph of the given system of equations. y + 2x = -1 3y - x = 4 graph of two lines that intersect at the point (negative 1, 1) graph of two lines that intersect at the point (1, 1) graph of two parallel lines with positive slopes None of the above I'm not understanding this one

OpenStudy (anonymous):

@ankitshaw

OpenStudy (ankitshaw):

where are the graphs? to choose

OpenStudy (anonymous):

Choose the correct graph of the given system of equations. y + 2x = -1 3y - x = 4 graph of two lines that intersect at the point (negative 1, 1) graph of two lines that intersect at the point (1, 1) graph of two parallel lines with positive slopes None of the above

OpenStudy (anonymous):

The graphs are those 4 choices

OpenStudy (ankitshaw):

okk now i understood

OpenStudy (ankitshaw):

 option 1 is correct

OpenStudy (ankitshaw):

plz once try it yourself after understanding

OpenStudy (anonymous):

Thank you. How did you get 8?

OpenStudy (ankitshaw):

when we distribute 2 in (3y-4) we get 2X3y - 2X4 = 6y -8